A 35-L block of ice (density at 0.9167 g/cm3) sits on a surface that is 25 inches wide and 16 inches long on Mars, where the acceleration due to gravity is only 38.74% that of Earth’s. Express the pressure exerted on this surface using atmospheres.


1
Expert's answer
2021-03-17T08:46:58-0400

m = "\frac{V}{ρ}" = 35000 / 0.9167 = 38180.43 g = 38.18043 kg

S = 25"\times"16 = 400 in2 = 0.258 m2

P = "\frac{mg}{S}" = "\frac{38.18043\times(0.3874\times9.81)}{0.258}" = 562.4 Pa = 0.00555 atm.